# On the Relationship between Space-Time Accessibility and Leisure Activity Participation

### v1

@@ 5.2 Selectivity in spatial choice of leisure activities
Based on 856 individuals where their $A_{i}>0$ and have leisure activities observed, we show a strong alignment between SPAi and actual choices of leisure activities (Figure 7). Only 0.1% of individuals have zero visited places within their capability set, the distribution of visited places outside SPAi is skewed toward lower values (Figure 7a): with a bottom 10% who are mainly car users having over 50% of their visits outside their capability set. Individuals who frequently travel beyond their SPA are more likely to rely on cars. Space-time accessibility in this context primarily captures trip-chaining opportunities between home and work, framing leisure visits as secondary or incidental. While informative, this representation can be inadequate for modeling car users who travel for longer than 90 minutes, whose greater spatial flexibility allows for more diverse and decoupled travel patterns.
@@
Most individuals exhibit negative $d_{i}$ values (Figure 7b), indicating a tendency to choose better-ranked locations than expected by chance, as 67% of individuals have significant deviation from being random ($p<0.05$). This supports the feasibility of using SPA as a realistic approximation of individuals’ capability sets. A cluster of individuals shows strongly negative $d_{i}$ with highly significant $p$-values, while others have values near zero or even positive, suggesting closer-to-random or worse-than-random behavior. These results reveal considerable heterogeneity in the degree of behavioral selectivity across the population.
@@
Among all individual attributes considered, only the use of active transport mode shows a statistically significant association with the effect size ($d_{i}$), as illustrated in Figure 7c. Non-active users exhibit a higher median $d_{i}$ and a tighter distribution skewed toward weaker selectivity (i.e., more positive values). In contrast, active mode users show stronger selectivity as expected by their capability set ($\beta=-0.51,p<0.001$). These results suggest that individuals using active mode tend to align more closely with their SPA-based opportunity rankings, whereas car users may be less constrained or freer in their spatial choices.
@@
Individuals exhibit varying degrees of alignment with modeled opportunity structures. Car users have a higher chance of traveling beyond their capabilities. Meanwhile, active users show stronger within-SPA selectivity. These findings underscore the value of decomposing behavioral mechanisms underlying accessibility: individuals do not uniformly act upon modeled opportunities, and their selectivity depends critically on their transport mode and spatial constraints.
@@
Figure 7: Selectivity in spatial choice behavior across individuals. a, Distribution of individuals by the share of visited locations outside their modeled feasible set (SPAi), with bar color indicating the proportion of car users in each bin. b, Standardized effect size $d_{i}$ plotted against the empirical $p$-value from the mean rank test, on a log scale. The vertical line marks the $p=0.05$ threshold. Each point is an individual. c, Boxplot of $d_{i}$ stratified by use of active transport mode, with jittered points colored by statistical significance (blue: $p<0.05$).
@@ 5.3 Modeling outcomes
We apply structural equation modeling techniques with the diagonally weighted least squares (DWLS) estimator on a total of 2,415 weighted observations (individuals). The model demonstrated excellent overall fit, with a Comparative Fit Index (CFI) of 0.998 and Tucker–Lewis Index (TLI) of 0.963. The Root Mean Square Error of Approximation (RMSEA) was 0.041, with a 90% confidence interval of [0.018, 0.067], and the Standardized Root Mean Square Residual (SRMR) was 0.008, all supporting model adequacy (Hu and Bentler, 1999).
Significant direct and indirect pathways with standardized coefficients exceeding an absolute value of 0.1 (Figure 8) highlight the role of household structure, active mode use, and public transport subscription in shaping space–time accessibility, trip making, and leisure activity outcomes.
Figure 8: Structural equation model of factors shaping leisure activity participation. Standardized path coefficients ($p<0.05$) are shown along the arrows, with positive effects in teal and negative effects in orange. Individual attributes (left) influence transport mode choices and space–time accessibility (center), which in turn shape travel behavior (trip making) and leisure activity participation (right).
@@ 5.3.1 Transport mode and space-time accessibility
Household structure and lifestyle indicators emerged as strong predictors of transport mode choice. In particular, living in a couple with children substantially increases the likelihood of using the car as the main mode ($\beta=0.17$, $p<.001$), which aligns with prior evidence that households with children face greater time constraints and logistical needs, thereby relying more heavily on private vehicles (Scheiner and Holz-Rau, 2013). In contrast, individuals reporting active mode use are much less likely to depend on the car ($\beta=-0.63$, $p<.001$).
@@
Public transport subscription also shows important interactions with lifestyle and household type. While subscription is positively associated with active mode use ($\beta=0.11$, $p<.001$), reflecting complementarities between walking/cycling and transit (Oeschger et al., 2020), it is negatively associated with family-related household types, such as couples with children ($\beta=-0.19$, $p<.001$). This suggests that households with children are less able to rely on public transit, possibly due to constraints of escorting children and multi-stop trip chaining (Scheiner and Holz-Rau, 2013).
@@
Space-time accessibility is shaped by both sociodemographic and transport characteristics. Active mode users—those regularly walking or cycling—exhibit significantly higher SPA ($\beta=0.17$, $p<.001$). Given how SPA is evaluated, this could be ascribed to two main mechanisms: (1) these individuals mostly live in the city of Paris, with good proximity to leisure opportunities and good support for walking and cycling; and (2) they have relatively short commuting times (34 min) compared to non-active mode user (43 min), making the assumed 90-minute daily travel time budget have a greater share available for potential leisure activities.
@@
The positive relationship between car usage and STA ($\beta=0.29$, $p<.001$), while public transport subscription is negatively associated with STA ($\beta=-0.18$, $p<.001$). This suggests that, given individuals’ fixed daily anchors (home and work/study locations) and their chosen mode of travel, the car provides more effective access to a wider range of leisure opportunities compared to public transport.
@@
Household structure itself significantly shapes accessibility. Respondents living in couples with children yield lower STA ($\beta=-0.1$, $p=.001$). One explanation is that caregiving responsibilities often reduce spatial and temporal flexibility, leading to a tight travel time budget for planned activities.
5.3.2 Trip making
@@
Space-time accessibility strongly predicts total travel time ($\beta=-0.37$, $p<.001$). Good SPA implies a greater number of accessible leisure opportunities within the time–space prism, which in turn leads to shorter travel times, indicating that part of the activity demand can be met in proximity. The use of active modes is also associated with shorter total travel time, suggesting that those who rely on active transport do so because their activity needs can be met closer to their home and work anchors. The total travel time, in turn, serves as a mediator between SPA and leisure activity participation.
@@ 5.3.3 Leisure activity participation
The leisure activity diversity is shaped by space-time accessibility, travel behaviors, and individual attributes. Higher accessibility is positively associated with greater location diversity ($\beta=0.13$, $p<.001$). This finding underscores a conceptual alignment between space-time accessibility to leisure opportunities and actual participation in diverse leisure activities. In other words, individuals with higher accessibility to leisure opportunities are more likely to translate potential into visits to a broader range of leisure locations.
Trip making also contributes here: longer travel times increase leisure activity time ($\beta=0.23$, $p=<.001$). From the perspective of utility-maximization, longer travel time can still result in higher overall utility if it enables meaningful leisure engagement (De Vos et al., 2016). This effect is even stronger than that of SPA directly, underscoring the role of total travel time as a key mediator shaping actual leisure activity participation.
Lifestyle differences are notable. Active mode use is positively associated with leisure activity location diversity ($\beta=0.23$, $p<.001$). Leisure travel is often motivated by social needs, the pursuit of variety, and personal recharge (Stauffacher et al., 2005). Active transport may facilitate these motivations by providing greater flexibility in route selection and opening up opportunities to access varied leisure spaces such as parks, cafés, or community centers. It’s plausible that walking or cycling itself becomes part of a user’s leisure experience—adding intentionality and value to the travel act.
@@ 5.3.4 Effect decompositions of space-time accessibility
Decomposition analyses show that space–time accessibility had both direct and indirect effects on leisure participation (Table 2). The direct effect of accessibility on leisure diversity is positive and significant ($\beta=0.13$, $p<.001$), while the indirect effect through travel time is negative and significant ($\beta=-0.09$, $p<.001$). As a result, the total effect is small and marginally statistically significant ($\beta=0.04$, $p=0.077$), indicating that positive direct influences are offset by negative indirect pathways. In addition, active mode use exerted strong positive direct effects on leisure participation ($\beta=0.24$, $p<.001$), underscoring heterogeneity in activity outcomes beyond accessibility constraints.
@@
Direct Indirect Total
Effect on total travel time
   SPA ($A_{i}$) $^{***} 0.37-0.37^{***}$
@@
Effect on leisure diversity
   SPA ($A_{i}$) $13^{***} 0.0$$8^{***} 0.0$40.04
   Travel time $^{***} 0.230.23^{***}$
@@
Notes: SPA = space–time accessibility. Indirect effect of SPA operates through total travel time. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
While the total effect of SPA on leisure participation was not strongly significant ($p=0.077$), the decomposition highlights that SPA does matter: it exerts a positive direct effect on activity diversity, but this is offset by a negative indirect effect operating through reduced travel time. These countervailing mechanisms suggest that accessibility shapes leisure participation in complex ways, rather than through a simple net effect.
@@ 6 Conclusion
This study investigated how individual-based space–time accessibility (SPA) influences leisure activity participation, with a focus on the Paris region. By leveraging high-resolution mobility data, multimodal transport networks, and a capability-based accessibility framework, we demonstrated that SPA—operationalized as the feasible set of leisure opportunities within individuals’ time budgets—plays a critical role in shaping both the efficiency of travel (via total travel time) and the diversity of leisure activity locations.
Our findings provide strong empirical support for the hypothesis that higher SPA enables more efficient traveling in meeting daily activity demand and richer leisure activity engagement. Through structural equation modeling, we identified both direct and indirect pathways: SPA directly promotes greater leisure diversity, but it also reduces travel time, which is associated with lower diversity, resulting in offsetting effects. The observed behavioral selectivity in spatial choices further validated the SPA construct as a realistic proxy of individuals’ capability sets—particularly among active transport users, who showed stronger alignment with modeled opportunity structures.
Importantly, the study revealed substantial heterogeneity in accessibility and activity outcomes across population groups. Household structures involving caregiving responsibilities (e.g., single parents, couples with children) display lower SPA and reduced leisure participation, pointing to persistent structural constraints. These findings underscore the importance of considering both the spatial distribution of opportunities and the temporal feasibility of access when designing inclusive transport policies.
By embedding accessibility into a human capabilities framework, this work shifts the discourse from infrastructure provision toward understanding how individuals convert transport resources into meaningful participation. Future research should extend this analysis to non-commuting populations, incorporate dynamic time budgets, and examine the longitudinal impacts of accessibility improvements on social inclusion and well-being. As cities pursue more equitable and sustainable mobility systems, metrics like space-time accessibility can serve as valuable tools to identify gaps, monitor policy impacts, and prioritize interventions that expand real freedoms to engage in everyday urban life.
@@
The data used in this study were provided through participation in the NetMob 2025 Data Challenge (Chasse et al., 2025), under a non-disclosure agreement (NDA) between all authors, Inria with IFPEN. Due to licensing terms and privacy constraints governed by the European General Data Protection Regulation (GDPR), access to the data is restricted. Venue locations and categories can be retrieved from Overture API. Census data (income) were collected from Institut national de la statistique et des études économiques (INSEE) that is publicly available. GTFS data were collected from transport.data.gouv.fr that are publicly available. All data were utilized in accordance with the terms of service specified by their respective provider.

### v2

@@ 5.2 Selectivity in spatial choice of leisure activities
Based on 856 individuals where their $A_{i}>0$ and have leisure activities registered, we observe a strong alignment between STAi and actual spatial choices of leisure activities (Figure 8). The distribution of visited locations outside STAi is skewed toward lower values (Figure 8a): the top 10% of individuals make more than half of their visits outside their capability set. Moreover, only 0.1% of individuals have visited zero places within their capability set. Space-time accessibility in this context primarily captures trip-chaining opportunities between home and work, framing leisure visits as secondary or incidental.
@@
Most individuals exhibit negative $d_{i}$ values (Figure 8b), indicating a tendency to choose closer locations than expected by chance, as 67% of individuals have significant deviation from being random ($p<0.05$). This supports the feasibility of using STA as a realistic approximation of individuals’ capability sets. A cluster of individuals exhibits strongly negative $d_{i}$ values with highly significant $p$-values, while others have values near zero or even positive, suggesting behavior closer to random or worse than random. These results reveal considerable heterogeneity in the degree of behavioral selectivity across the population.
@@
Among all individual attributes considered, only the use of active transport mode shows a statistically significant association with the effect size ($d_{i}$), as illustrated in Figure 8c. Active users exhibit a higher median $d_{i}$ and a tighter distribution skewed toward weaker selectivity (i.e., more positive values). Individuals who use active modes exhibit significantly higher $d_{i}$ than non-active travelers, with an estimated increase of 0.27 units ($p=0.013$), suggested by the linear regression test $d_{i}\sim C(\text{Using Active Mode})$. These results suggest that individuals using active mode tend to align more closely with their STA-based opportunity rankings, whereas car users may be less constrained or freer in their spatial choices.
@@
Individuals exhibit varying degrees of alignment with modeled opportunity structures. Active transport users show stronger within-STA selectivity. These findings underscore the value of decomposing the behavioral mechanisms underlying accessibility: individuals do not uniformly act upon modeled opportunities, and their selectivity depends critically on their lifestyle and spatial constraints.
@@
Figure 8: Selectivity in spatial choice behavior across individuals. a, Distribution of individuals by the share of visited locations outside their modeled feasible set (STAi), with bar color indicating the proportion of car users in each bin. Bins include lower boundaries. b, Standardized effect size $d_{i}$ plotted against the empirical $p$-value from the mean rank test, on a log scale. The vertical line marks the $p=0.05$ threshold. Each point is an individual. c, Boxplot of $d_{i}$ stratified by use of active transport mode, with jittered points colored by statistical significance (blue: $p<0.05$).
@@ 5.3 Modeling outcomes
We apply structural equation modeling techniques with the diagonally weighted least squares (DWLS) estimator on a total of 2,415 weighted observations (individuals) to model leisure activity participation. The model demonstrated excellent overall fit, with a Comparative Fit Index (CFI) of 0.998 and Tucker–Lewis Index (TLI) of 0.963. The Root Mean Square Error of Approximation (RMSEA) was 0.041, with a 90% confidence interval of [0.018, 0.067], and the Standardized Root Mean Square Residual (SRMR) was 0.008, all supporting model adequacy (hu1999cutoff).
Significant direct and indirect pathways with standardized coefficients exceeding an absolute value of 0.1 (Figure 9) highlight the role of household structure, active mode use, and public transport subscription in shaping space–time accessibility, total travel time, and leisure activity outcomes.
Figure 9: Structural equation model of factors shaping leisure activity participation. Standardized path coefficients ($p<0.05$) are shown along the arrows, with positive effects in teal and negative effects in orange. Individual attributes (left) influence transport mode choices and space–time accessibility (center), which in turn shape travel behavior and leisure activity participation (right).
@@ 5.3.1 Transport mode and space-time accessibility
Household structure and lifestyle indicators emerged as strong predictors of transport mode choice. In particular, living in a couple with children substantially increases the likelihood of using the car as the main mode ($\beta=0.17$, $p<.001$), which aligns with prior evidence that households with children face greater time constraints and logistical needs, thereby relying more heavily on private vehicles (scheiner2013comprehensive). In contrast, individuals reporting active mode use are much less likely to depend on the car ($\beta=-0.63$, $p<.001$).
@@
Public transport subscription also shows important interactions with lifestyle and household type. While subscription is positively associated with active mode use ($\beta=0.11$, $p<.001$), reflecting complementarities between walking/cycling and transit (oeschger2020micromobility), it is negatively associated with family-related household types, such as couples with children ($\beta=-0.19$, $p<.001$). This suggests that households with children are less able to rely on public transit, possibly due to constraints of escorting children and multi-stop trip chaining (scheiner2013comprehensive).
@@
Space-time accessibility is shaped by both sociodemographic and transport characteristics. Active mode users—those regularly walking or cycling—exhibit significantly higher STA value ($\beta=0.17$, $p<.001$). Given how STA is evaluated, this could be ascribed to two main mechanisms: (1) these individuals mostly live in the city of Paris, with good proximity to leisure opportunities and good support for walking and cycling; and (2) they have relatively short commuting times (34 min) compared to non-active mode user (43 min), making the assumed 90-minute daily travel time budget have a greater share available for potential leisure activities.
@@
We find car usage to be positively associated with STA ($\beta=0.29$, $p<.001$), while public transport subscription is negatively associated with STA ($\beta=-0.18$, $p<.001$). This suggests that, given individuals’ fixed daily anchors (home and work/study locations) and their chosen mode of travel, the car provides more effective access to a wider range of leisure opportunities compared to public transport.
@@
Household structure itself also significantly shapes accessibility. Respondents living in couples with children yield lower STA ($\beta=-0.1$, $p=.001$). One explanation is that caregiving responsibilities often reduce spatial and temporal flexibility, leading to a tight travel time budget for planned activities (schwanen2008exploring).
5.3.2 Mobility
@@
Space-time accessibility strongly predicts total travel time ($\beta=-0.37$, $p<.001$). Good STA implies a greater number of accessible leisure opportunities within the time–space prism, which in turn leads to shorter travel times, indicating that part of the activity demand can be met in proximity. The use of active modes is also associated with shorter total travel time, suggesting that those who rely on active transport do so because their activity needs can be met closer to their home and work anchors. The total travel time, in turn, serves as a mediator between STA and leisure activity participation.
@@ 5.3.3 Leisure activity participation
The leisure location diversity is shaped by space-time accessibility, travel behaviors, and individual attributes. Higher accessibility is positively associated with greater location diversity ($\beta=0.13$, $p<.001$). This finding underscores a conceptual alignment between space-time accessibility to leisure opportunities and actual participation in diverse leisure locations. In other words, individuals with higher accessibility to leisure opportunities are more likely to translate potential into visits to a broader range of leisure locations.
Mobility also contributes here: longer travel times increase leisure location diversity ($\beta=0.23$, $p=<.001$). From the perspective of utility-maximization, longer travel time can still result in higher overall utility if it enables meaningful leisure engagement (de2016travel). This effect is even stronger than that of STA directly, underscoring the role of total travel time as a key mediator shaping actual leisure activity participation.
Lifestyle differences are notable. Active mode use is positively associated with leisure activity location diversity ($\beta=0.23$, $p<.001$). Leisure travel is often motivated by social needs, the pursuit of variety, and personal recharge (stauffacher2005diversity). Active transport may facilitate these motivations by providing greater flexibility in route selection and opening up opportunities to access varied leisure spaces such as parks, cafés, or community centers. It’s plausible that walking or cycling itself becomes part of a user’s leisure experience—adding intentionality and value to the travel act.
@@ 5.3.4 Effect decompositions of space-time accessibility
Decomposition analyses reveal that the space-time accessibility value had both direct and indirect effects on leisure participation (Table 2). The direct effect of accessibility on leisure location diversity is positive and significant ($\beta=0.13$, $p<.001$), while the indirect effect through travel time is negative and significant ($\beta=-0.09$, $p<.001$). As a result, the total effect is small and marginally statistically significant ($\beta=0.04$, $p=0.077$), indicating that positive direct influences are offset by negative indirect pathways. In addition, active mode use exerted a strong, positive direct effect on leisure participation ($\beta=0.24$, $p<.001$), underscoring heterogeneity in activity outcomes beyond accessibility constraints.
@@
Effect on Direct Indirect Total
Total travel time
   STA value ($A_{i}$) $^{***} 0.37-0.37^{***}$
@@
Leisure location diversity
   STA value ($A_{i}$) $13^{***} 0.0$$8^{***} 0.0$40.04
   Total travel time $^{***} 0.230.23^{***}$
@@
Notes: STA = space–time accessibility. The indirect effect of STA operates through the total travel time. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
While the total effect of STA on leisure participation was not strongly significant ($p=0.077$), the decomposition highlights that STA does matter: it exerts a positive direct effect on leisure location diversity, but this is offset by a negative indirect effect operating through reduced travel time. These countervailing mechanisms suggest that accessibility shapes leisure activity participation in complex ways, rather than through a simple net effect.
@@ 6 Conclusion
This study examined the impact of individual-based space–time accessibility (STA) on leisure activity participation, with a focus on the Paris region. By leveraging high-resolution GPS data for 2,415 Paris residents, multimodal transport networks, and a capability-based accessibility framework, we demonstrated that STA—operationalized as the feasible set of leisure opportunities within individuals’ time budgets—plays a critical role in shaping both the efficiency of travel (via total travel time) and the diversity of leisure activity locations.
Our results offer robust empirical evidence that greater transport accessibility promotes more efficient travel (i.e., shorter travel times) to fulfill daily activity demand and supports broader participation in leisure activities (i.e., location diversity). Through structural equation modeling, we identified both direct and indirect effects: STA directly promotes a greater leisure location diversity, but it also reduces travel time, which is associated with lower diversity, resulting in offsetting effects. The observed spatial choices of leisure destinations further validated the STA construct as a realistic proxy for individuals’ capability sets—particularly among active transport users, who showed a stronger alignment with modeled opportunity structures.
Importantly, the study revealed substantial heterogeneity in accessibility and activity outcomes across population groups. Household structures involving caregiving responsibilities (e.g., single parents, couples with children) exhibit lower STA and reduced leisure participation, indicating persistent structural constraints. These findings underscore the importance of considering both the spatial distribution of opportunities and the temporal feasibility of access when designing inclusive transport policies.
By embedding accessibility into a human capabilities framework, this work shifts the discourse from infrastructure provision toward understanding how individuals convert transport resources into meaningful participation. Future research should extend this analysis to non-commuting populations, incorporate dynamic time budgets, and examine the longitudinal impacts of accessibility improvements on social inclusion and well-being. As cities pursue more equitable and sustainable mobility systems, more nuanced accessibility metrics, such as space-time accessibility, can serve as valuable tools to identify gaps, monitor policy impacts, and prioritize interventions that expand real freedoms to engage in everyday urban life.
@@
The data used in this study were provided through participation in the NetMob 2025 Data Challenge (netmob25), under a non-disclosure agreement (NDA) between all authors, Inria with IFPEN. Due to licensing terms and privacy constraints governed by the European General Data Protection Regulation (GDPR), access to the data is restricted. Venue locations and categories can be retrieved from Overture API. Census data (income) were collected from Institut national de la statistique et des études économiques (INSEE) that is publicly available. GTFS data were collected from transport.data.gouv.fr that are publicly available. All data were utilized in accordance with the terms of service specified by their respective provider.